How can Al^3+ and S^2- be packed in a unit cell to make a neutral compound? The answer is 'Al on the corners and in center, S on edges', but I don't understand why.
Answer #1

To make a neutral compund charge on cation and anion should be same.

Atom at corner contribution = 1/8

Atom at center = 1

Atom at edge centre = 1/4

So total aluminum atom = corner +. Center = 8*1/8 + 1 = 2 atom

Total charge on Al = +3*2 = +6

Total Sulfur atom = edge centre = 12 *1/4 = 3

Total charge on S - atom = -2 *3 = +6

So neutral compund is formed in unit cell.